OpenStudy (anonymous):

Solve and check. Label any contradictions or identities. what did i do wrong? – 8a/7 – 2 = 4 -8a/7*7/8=4+2 a=6 Check: 8*6/7= 4 48/7=4 6 6/7-2 =4 4 6/7= 4 Thus -8a/7-2=4 is in fact an Contradication.

OpenStudy (anonymous):

no, you did it incorrectly. -8a/7 = 6 now multiply both sides by 7/8 you forgot to multiply on the RHS

OpenStudy (anonymous):

OK i didn't multiple the other side that's why is wrong

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

8a/7= 6 8/7a / 7/8 = 6/1 * 8/6 a = 48/6 a=8 Thus -8a/7-2=4 is in fact an Contradication.

OpenStudy (anonymous):

why are you multiplying the right hand side by 8/6?

OpenStudy (anonymous):

sorry made a mistake. i meant this 8a/7= 6 8/7a / 7/8 = 6/1 * 7/8 a = 36/8 a= 4 1/2 Thus -8a/7-2=4 is in fact an Contradication.

OpenStudy (anonymous):

6 times 7 is 42. Please do your calculations properly.

OpenStudy (anonymous):

you are correct i should of used the calculator. thanks for the correction

OpenStudy (anonymous):

so the answer is 5.25 Or 5 1/4

OpenStudy (anonymous):

if you are solving an equation, you cannot get a contradiction, because when you solve an equation for a variable, that variable is what solves the equation. In other words, that variable is the missing link.

OpenStudy (anonymous):

So its is an identiy. this is what i need to answer weather it's an label of any contradictions or identities?

OpenStudy (anonymous):

But it's not equal to? I thought when u solve the equation and one side is not equal to the other side then that's an contradiction?

OpenStudy (anonymous):

OK understood.

OpenStudy (anonymous):

no, when you solve an equation, you are finding out the value that "solves the equation". for example, x +3 = 5 subtract 3 from both sides x = 2 I solved the equation x+3 = 5 and got x to be the value 2 that means this value of x cannot contradict the equation because I solved the equation to get the value of x as 2.

OpenStudy (anonymous):

the very fact that I solved the equation means that the value I got as the missing link cannot be a contradiction. do you get it?

OpenStudy (anonymous):

I understand that but when i found the answer for a in my problem it didn't equal 4 which is what it was suppose to do. I don't know how to get it to equal to 4 without changing the entire problem

OpenStudy (anonymous):

if you solve correctly, the value you find for a will ALWAYS solve the problem. It just means you made some error in calculating the value.

OpenStudy (anonymous):

-8a/7 - 2 = 4 add 2 to both sides -8a/7 = 6 a = -42/8 a = -5.25 now, substitute -5.25 in the original equation -8(-5.25)/7 - 2 = 4 LHS = -8(-5.25)/7 -2 = 42/7 -2 = 6-2 = 4 = RHS

OpenStudy (anonymous):

exactly and i believe it's in what i am multiplying. i have to be more careful when I'm typing the numbers because that where my error are coming from. thanks so much for the help.

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

make sure you do the calculations correctly.